OpenStudy (anonymous):

During a weekend sale, various articles were reduced in price by the indicated percent. By what percent would the price of each one have to be increased after the sale to return to the original price? a. ladies coats reduced 30 percent b. ovens reduced 25 percent c. blue jeans reduced 40 percent d. mountain bikes reduced 15 percent

OpenStudy (anonymous):

a) ok, let's say ladies coats cost 100$ a piece. we will give the article a example price to make it easier to understand the assignment assigment says they were reduced by 30%, that means: their price on sale was 100-30 = 70% of their original price 70% * 100$ = 0.7*100 = 70$ now the price has to return from 70$ (sale price) to 100$ (original price). that actually means they ask you how much percent of 70$ is 30$, or formula: 70*x=30 x = 30/70 = 0.428 *100% = 42.8% now you can try the same thing with b,c,d examples. feel free to ask if you get stuck;)

OpenStudy (anonymous):

thanks a lot! ive gotta be going out, so I hope to talk to you again! thank you for all your help and the formulas!:) see ya! thanks so much!!!